OpenStudy (elite):

The solution to 4x + 7 - 9 = -14 is a0.

alones (alones):

Let's solve your equation step-by-step. 4x+7−9=−14 Step 1: Simplify both sides of the equation. 4x+7−9=−14 Simplify: ( steps if u need 'em) 4x+7+−9=−14 (4x)+(7+−9)=−14(Combine Like Terms) 4x+−2=−14 4x−2=−14 Step 2: Add 2 to both sides. 4x−2+2=−14+2 4x=−12 Step 3: Divide both sides by 4. 4x4=−124 x=−3 hope that helps =]

OpenStudy (elite):

ok thanks step by step is better

alones (alones):

IKR :P